Solve: Log2(2) +log2(x+3)=5
Answer by stanbon(75887)

Solve: Log2(2) +log2(x+3)=5
--------------
Log2[2(x+3)) = 5
2(x+3) = 2^5
2x + 4 = 32
2x = 28
x = 14
--------------
Cheers,
Stan H.
------------